# A Multi-Level Superoptimizer for Tensor Programs

### v2

@@ 8.1. Experimental Setup
8.2. Benchmark Results
Figure 7 compares the performance of Mirage with existing tensor program optimizers on six DNN benchmarks on NVIDIA A100 and H100 GPUs. All systems use half-precision floating points to serve these DNN benchmarks. TASO (TASO) and PET (wang2021pet) are DNN superoptimizers that automatically generate graph optimizations at the kernel level. PyTorch (pytorch) uses the highly engineered cuDNN and cuBLAS libraries (cublas; cudnn) to perform DNN operators on GPUs. TensorRT and its LLM variant TensorRT-LLM include a set of manually designed and highly optimized kernels for common tensor operators such as attention (tensorrt). FlashAttention and its inference variant FlashDecoding are manually designed and implemented kernels for fast attention computation (dao2023flash; hong2024flashdecoding). Finally, Triton (tillet2019triton) is a schedule-based optimizer to generate high-performance kernels and has been deployed in existing systems, achieving better performance than other schedule-based optimizers.
Compared to the best existing approaches, Mirage improves the performance of these benchmarks by 1.1-2.9$\times$ by combing algebraic transformations, schedule transformations, and the generation of new custom kernels. §3 shows the best discovered $\mu$Graphs for RMSNorm. Next, we present a case study for the other benchmarks.
@@ 8.2. Benchmark Results
Figure 8. Comparing the $\mu$Graphs used by existing optimizers and Mirage for QKNorm and attention.
@@
Group-query attention is the backbone of LLMs and has been heavily optimized by existing frameworks. For example, FlashAttention and FlashDecoding are expert-designed attention kernels and have been adopted in existing LLM inference systems (dao2023flash). Mirage discovers existing expert-designed kernels as well as other $\mu$Graphs that outperform them by up to 2.2$\times$. The speedup is achieved by two additional optimizations on top of existing hand-written kernels. First, current approaches rely on heuristics to determine the grid dimensions for GQA, which are suboptimal in certain scenarios. For example, TensorRT-LLM launches the attention kernel with grid dimensions of (8, 2, 1) and (8, 2, 8) for GQA when the batch sizes are 1 and 8, respectively. Both configurations cannot fully utilize all A100 and H100 GPUs, which have 108 and 132 SMs respectively. On the other hand, Mirage automatically searches for the best grid dimensions for each $\mu$Graph, resulting in fully utilized SMs.
Second, existing approaches use fixed tensor dimensions to parallelize GQA across thread blocks. For example, FlashAttention (dao2023flash) parallelizes attention across blocks using the sample, head, and query sequence dimensions, while FlashDecoding and TensorRT-LLM leverage the sample, head, and key-value sequence dimensions. Both strategies are efficient for conventional multi-head attention with many attention heads but suboptimal for GQA with limited attention heads. In contrast, Mirage automatically selects the most efficient dimensions among sample, KV heads, query sequence, and key-value sequence, and uses different $\mu$Graphs for different attention scenarios, reducing device memory access by up to 7$\times$ compared to the heuristics used in existing systems. Implementing Mirage’s $\mu$Graphs in existing systems is possible but requires extensive engineering effort to support different kernels for different scenarios, while Mirage automatically generates them and verify their correctness.
@@
To reduce model divergence, several recent DNNs introduce query-key normalization (QKNorm) into the Transformer architecture (chameleonteam2024chameleon). QKNorm applies layer normalization to the query and key vectors before attention, as shown in Figure 8a. These additional normalization layers are not yet supported by existing attention implementations (e.g., FlashAttention and TensorRT-LLM) and require launching separate kernels for normalization and attention.
Mirage automatically discovers a $\mu$Graph that integrates QKNorm and attention computation into a custom kernel, as shown in Figure 8b. The $\mu$Graph reorganizes the attention computation to enable fusion with the two layer normalizations, which avoids writing intermediate results to GPU device memory and reduces the kernel execution time by 1.1-1.4$\times$.
@@
Figure 9. Comparing the tensor programs used by existing optimizers and by Mirage for LoRA: $O=W\times X+B\times A\times X$. Note that both matrices $A$ and $B$ are low-rank.
@@
Low-rank adaptation (LoRA) introduces a pair of low-rank adapters to the linear operators of a pre-trained DNN to improve its predictive performance for downstream tasks. Existing tensor program optimizers launch separate kernels for the original linear operator and the two new linear operators in LoRA (Figure 9a), which introduces high kernel launch overheads since the LoRA operators involve very low computational costs. Figure 9b shows the best $\mu$Graph discovered by Mirage for LoRA, which fuses the three Matmuls and the subsequent Add into a single kernel. Mirage reorganizes the computation into two thread-block level Matmuls by leveraging the following algebraic transformation: $W\times X+B\times A\times X=(W\|B)\times\big{(}X\|(A\times X)\big{)}$. The Concats in Figure 9b do not involve any computation and are performed by updating tensor offsets in GPU shared memory. This $\mu$Graph reduces the execution cost of LoRA by 1.7-1.8$\times$.
@@
Figure 10. Comparing the $\mu$Graphs used by existing optimizers and Mirage for GatedMLP.
@@
Gated multi-layer perceptron is commonly used in DNNs to capture non-linear representations. We use the GatedMLP configuration introduced in Falcon-7B (falcon40b), whose kernel graph is shown in Figure 10a. Existing tensor program optimizers generally fuse the two Matmuls in a single kernel to reduce GPU device memory access since the input tensor $X$ only needs to be loaded once. However, this approach still requires launching multiple kernels and storing intermediate results (i.e., the output of the two Matmuls) in device memory as the SiLU activation and elementwise multiplication are not fused with the Matmuls.
In contrast, the best $\mu$Graph discovered by Mirage (Figure 10b) performs the two Matmuls in parallel in the same block graph and fuses the remaining computation (i.e., SiLU and Mul) as a post-processing step into the block graph. This approach yields 1.4-1.5$\times$ speedups on A100 GPUs and 2.7-2.9$\times$ speedups on H100 GPUs.
@@
To accelerate model training, nGPT introduces normalized Transformer, which normalizes all intermediate results in Transformer. Formally, the computation is defined as $y=\texttt{Norm}(x+\alpha(\texttt{Norm}(h-x)))$, where Norm is a normalization layer, and $x$, $h$, and $\alpha$ are the input tensors. Existing systems launch three separate kernels for nTrans since it interleaves normalization and elementwise addition and multiplication. Mirage automatically discovers a $\mu$Graph that fuses the computation into a single kernel and stores all intermediate results in the GPU shared memory, reducing the kernel execution time by 1.8-2.0$\times$ compared to the best existing systems.
Figure 11. Comparing the end-to-end inference performance of PyTorch and PyTorch with Mirage-generated kernels.
8.3. End-to-end Results
In addition to the microbenchmark performance, we also evaluate how Mirage-generated kernels reduce the end-to-end latency of commonly used DNNs. Mirage supports just-in-time compilation and deployment, and its generated kernels can be directly integrated in PyTorch programs. We compare PyTorch with its native handwritten CUDA kernels and PyTorch with Mirage-generated kernels on four DNN models. Figure 11 shows the results. Mirage reduces the end-to-end latency of these models by 1.1-1.9$\times$ by automatically generating highly optimized kernels. Note that the improvement is achieved by a few lines of code changes to the PyTorch programs.
8.4. Optimization Time and Ablation Study
Table 5. Ablation study on Mirage’s techniques to accelerate $\mu$Graph generation. We evaluate the impact of multi-threading and abstract expressions on search time for RMSNorm.
@@ 8.4. Optimization Time and Ablation Study
Generally, Mirage takes up to 4 hours to optimize a Lax program. Note that this optimization is a one-time cost before deployment on the target hardware. This section provides more detailed results and an ablation study for Mirage’s search procedure. We evaluate how our techniques allow Mirage to explore large $\mu$Graphs while maintaining low search time. We focus on two techniques: pruning via abstract expressions (§4.3) and multi-threading. Table 5 reports the search times for RMSNorm as we adjust the maximum number of operators in a block graph. Multi-threading significantly reduces the search time, while pruning via abstract expressions is crucial for the scalability of Mirage. Specifically, the pruning techniques allow Mirage to explore $\mu$Graphs whose block graph can each have at most 11 operators, while disabling abstract expressions restricts Mirage to consider up to 6 operators in a block graph in order to finish the search in 10 hours. Note that Mirage needs to consider 11 operators in a block graph in order to discover the $\mu$Graph shown in Figure 4 for RMSNorm.
9. Related Work
@@
Many existing frameworks, such as TensorFlow XLA (tensorflow_xla; Tensorflow), PyTorch (pytorch), and TensorRT (tensorrt), rely on GPU experts to manually design and implement kernels for ML operators. Recently, significant engineering effort has been dedicated to manually designing, implementing, and optimizing GPU kernels for commonly used DNNs (known as foundation models (bommasani2022opportunities)). For example, to optimize attention computation (transformers), recent work has introduced various kernels based on FlashAttention (dao2023flash; tri2023flashdecoding; hong2024flashdecoding; fastertransformer). Due to the increasing complexity of modern GPU architectures (e.g., tensor cores in A100s (Markidis2018tensorcores) and thread block clusters in H100s (nvidia-h100)), manually designed kernels may miss subtle optimizations that are hard to discover manually.
@@
Superoptimization was originally introduced to find optimal code for instruction sequences (massalin; stoke; peephole). Recent work has applied superoptimization techniques to tensor programs (TASO; wang2021pet; zheng2023einnet; tensat). All these attempts only consider algebraic transformations at the kernel level and cannot discover sophisticated optimizations that require jointly considering algebraic and schedule transformations at all of the kernel, block, and thread levels. Our evaluation shows that Mirage largely outperforms existing DNN superoptimizers, demonstrating the significance of multi-level joint optimizations.
@@
Recent work has introduced ML compilers that automatically optimize the schedule of kernel executions on GPUs. For example, TVM (tvm; tvm_auto_tuner), Ansor (ansor), and Triton (tillet2019triton), among others (zheng2020flextensor; hagedorn2023graphene; feng2022tensorir) are based on the idea of algorithm-schedule separation introduced in Halide and search for an optimized schedule to execute a given algorithm on GPUs. Schedule-based approaches rely on users to explicitly specify the algorithm for each kernel and their performance is limited to the provided algorithms.
@@
Welder (shi2023welder) and ASPEN (park2023aspen) introduce multi-level tile graphs that bear some similarity to Mirage’s $\mu$Graphs, as both representations follow the GPU memory hierarchy. However, prior work focuses on scheduling transformations, while Mirage covers these optimizations and also considers algebraic transformations and discovery of new custom kernels. Most optimizations shown in the paper are outside the scope of prior work.
10. Conclusion
This paper proposes Mirage, the first multi-level superoptimizer for tensor programs. Mirage introduces a hierarchy graph representation to specify a tensor program at the kernel, thread block, and thread levels of the GPU compute hierarchy, and uses a novel pruning technique based on abstraction to significantly reduce the search space Mirage needs to consider while providing a certain optimality guarantee. Mirage outperforms existing tensor program optimizers by up to 3.5$\times$ even for DNNs that are widely used and heavily optimized.

### v3

@@ 8.1 Experimental Setup
8.2 Benchmark Results
Figure 7 compares the performance of Mirage with systems on six DNN benchmarks on NVIDIA A100 and H100 GPUs. All systems use half-precision floating points to run these DNN benchmarks. TASO [25] and PET [46] are DNN superoptimizers that automatically generate algebraic transformations at the kernel level. We report a combined TASO/PET baseline, as the latest TASO implementation includes PET’s partially equivalent transformations as special substitutions. PyTorch [34] uses the highly optimized cuDNN and cuBLAS libraries [16, 15] to perform DNN operators on GPUs. For the PyTorch baseline, we enable torch.compile and use FlashAttention kernels to maximize performance. TensorRT and its LLM variant TensorRT-LLM include a set of manually designed and highly optimized kernels for common tensor operators such as attention [42]. FlashAttention and its inference variant FlashDecoding are manually written kernels for efficient attention [17, 21]. Finally, Triton is a schedule-based optimizer to generate high-performance kernels and has been adopted in production systems, outperforming other schedule-based approaches [43]. All baselines use CUDA Graphs to minimize kernel launch overhead.
Compared to the best existing approaches, Mirage improves the performance of these benchmarks by up to $\times3$ by combing algebraic transformations, schedule transformations, and the generation of new custom kernels. §3 shows the best discovered $\mu$Graphs for RMSNorm. Next, we present a case study for the remaining benchmarks.
@@ 8.2 Benchmark Results
Figure 8: Comparing the $\mu$Graphs used by existing optimizers and Mirage for QKNorm and attention.
@@
Group-query attention is the backbone of LLMs and has been heavily optimized by existing frameworks. For example, FlashAttention and FlashDecoding are expert-designed attention kernels and have been adopted in existing LLM inference systems [17]. Mirage discovers these expert-designed kernels as well as other $\mu$Graphs that outperform them by up to 2.2$\times$. The speedup is achieved by two additional optimizations on top of existing hand-written kernels. First, current approaches rely on fixed heuristics to determine the grid dimensions for GQA, which are suboptimal in certain scenarios. For example, TensorRT-LLM launches the GQA kernel with grid dimensions of (8, 2, 1) and (8, 2, 8) when the batch sizes are 1 and 8, respectively. However, both configurations cannot fully utilize all SMs on A100 (108 SMs) and H100 (132 SMs) GPUs. In contrast, Mirage automatically searches for the best grid dimensions for each $\mu$Graph, resulting in full SM utilization. Further ablation study shows that the performance of the best $\mu$Graph discovered by Mirage degrades by 18% when using the same grid dimensions as TensorRT-LLM.
Second, existing approaches use fixed tensor dimensions to parallelize GQA across thread blocks. For example, FlashAttention [17] parallelizes attention across the sample, head, and query sequence dimensions, while FlashDecoding and TensorRT-LLM leverage the sample, head, and key-value sequence dimensions. Both strategies are efficient for conventional multi-head attention with many heads but suboptimal for GQA with fewer attention heads. In contrast, Mirage automatically selects the most efficient parallelization strategy by choosing among the sample, KV heads, query sequence, and key-value sequence dimensions. Moreover, Mirage generates different $\mu$Graphs tailored to different attention scenarios, reducing device memory access by up to 7$\times$ compared to the heuristics used in existing systems.
Implementing Mirage’s $\mu$Graphs in existing systems is possible but requires extensive engineering effort to support different kernels for different scenarios. In contrast, Mirage automatically generates them and verify their correctness.
@@
To reduce model divergence, several recent DNNs introduce query-key normalization (QKNorm) into the Transformer architecture [40]. QKNorm applies layer normalization to the query and key vectors before attention, as shown in Figure 8a. These additional normalization layers are not yet supported by existing attention implementations (e.g., FlashAttention and TensorRT-LLM) and require launching separate kernels for normalization and attention.
Mirage automatically discovers a $\mu$Graph that integrates QKNorm and attention computation into a custom kernel, as shown in Figure 8b. The $\mu$Graph reorganizes the attention computation to enable fusion with the two layer normalizations, which avoids writing intermediate results to GPU device memory and reduces the kernel execution time by up to 1.4$\times$.
@@
Figure 9: Comparing the tensor programs used by existing optimizers and by Mirage for LoRA: $O=W\times X+B\times A\times X$. Note that both matrices $A$ and $B$ are low-rank.
@@
Low-rank adaptation (LoRA) introduces a pair of low-rank adapters to the linear operators of a pre-trained DNN to improve its performance for downstream tasks. Existing tensor program optimizers launch separate kernels for the original linear operator and the two additional linear operators introduced by LoRA (Figure 9a), which introduces high kernel launch overheads since these LoRA operators involve minimal computation. Figure 9b shows the best $\mu$Graph discovered by Mirage for LoRA, which fuses the three Matmuls and the subsequent Add into a single kernel. Mirage reorganizes the computation into two block-level Matmuls by leveraging the following algebraic transformation: $W\times X+B\times A\times X=(W\|B)\times\big{(}X\|(A\times X)\big{)}$. The Concats in Figure 9b do not involve any computation and are performed by updating tensor offsets in GPU shared memory. This $\mu$Graph reduces the execution cost of LoRA by 1.1-2.4$\times$.
@@
Figure 10: Comparing the $\mu$Graphs used by existing optimizers and Mirage for GatedMLP.
@@
Gated multi-layer perceptrons are commonly used in DNNs to capture non-linear representations. We use the GatedMLP configuration introduced in Falcon-7B [10], whose kernel graph is shown in Figure 10a. Existing tensor program optimizers generally fuse the two Matmuls in a single kernel to reduce GPU device memory access, since the input tensor $X$ only needs to be loaded once. However, this approach still requires launching multiple kernels and storing intermediate results—specifically, the output of the two Matmuls—in device memory, as the SiLU activation and elementwise multiplication are not fused with the Matmuls.
In contrast, the best $\mu$Graph discovered by Mirage (Figure 10b) performs the two Matmuls in parallel within the same block graph and fuses the remaining computation (i.e., SiLU and Mul) as post-processing steps within the same block graph. This approach yields 1.5$\times$ speedups on A100 GPUs and 2.7-3.3$\times$ speedups on H100 GPUs.
@@
To accelerate model training, nGPT introduces normalized Transformer, which normalizes all intermediate results in Transformer [28]. Formally, the computation is defined as $y=\texttt{Norm}(x+\alpha(\texttt{Norm}(h-x)))$, where Norm is a normalization layer, and $x$, $h$, and $\alpha$ are input tensors. Existing systems launch three separate kernels for nTrans, since it interleaves normalization and elementwise addition and multiplication. Mirage automatically discovers a $\mu$Graph that fuses the computation into a single kernel and stores all intermediate results in GPU shared memory. Mirage outperforms other baselines but is slower than TensorRT. This performance gap is because Mirage loads data from global memory to shared memory and writes it back for each tensor in graph-defined kernels. This design improves memory efficiency and enables asynchronous pipelines. However, for kernels with light computation, the overhead of these memory transfers can dominate the kernel runtime. To mitigate this overhead, we plan to extend Mirage to support bypassing shared memory during data loading, therefore avoiding unnecessary data movement.
Figure 11: Comparing the end-to-end inference performance of PyTorch and PyTorch with Mirage-generated kernels.
8.3 End-to-end Results
In addition to the microbenchmark performance, we also evaluate how Mirage-generated kernels impact the end-to-end latency of commonly used DNNs. Mirage supports just-in-time compilation and deployment, and its generated kernels can be directly integrated into PyTorch programs. We compare PyTorch with its native handwritten CUDA kernels and PyTorch with Mirage-generated kernels on four DNN models. Figure 11 shows the results. Mirage reduces the end-to-end latency of these models by 0.9-1.9$\times$ by automatically generating highly optimized kernels. The improvement is achieved with a few lines of code changes to the PyTorch programs.
8.4 Search Time
Table 5: Ablation study on Mirage’s techniques to accelerate $\mu$Graph generation. We evaluate the impact of multi-threading and abstract expressions on search time for RMSNorm.
@@ 8.4 Search Time
In our evaluation, Mirage takes up to 4 hours to optimize a Lax program. This optimization is a one-time cost before deployment on the target hardware. This subsection provides detailed results and an ablation study of Mirage’s search procedure, focusing on how its techniques enable the exploration of large $\mu$Graphs while maintaining low search time. In particular, we evaluate the impact of two techniques: pruning via abstract expressions (§4.3) and multi-threading. Table 5 reports the search times for RMSNorm as we vary the maximum number of operators allowed in a block graph.
Multi-threading significantly reduces the search time, while pruning via abstract expressions is crucial for the scalability of Mirage. Specifically, the pruning techniques allow Mirage to explore $\mu$Graphs whose block graphs can each have at most 11 operators, while disabling abstract expression pruning restricts Mirage to handle block graphs with up to 6 operators within a 10-hour search window. Note that discovering the optimized $\mu$Graph for RMSNorm shown in Figure 3 requires exploring block graphs with 11 operators.
8.5 Ablation Study on Optimizations
Figure 12: Ablation study on optimizations used in Mirage. We evaluate the performance degradation when disabling each optimization independently. The evaluation is performed on A100 for GQA with batch size $1$.
We conduct an ablation study to evaluate the impact of thread graph construction and optimizations introduced in §6, including layout optimization, operator scheduling, and memory planning. Specifically, we measure the performance degradation of the best $\mu$Graph discovered by Mirage when each optimization is disabled independently. The study is conducted on an A100 using the GQA benchmark with a batch size of $1$. The results, shown in Figure 12, indicate that disabling any individual optimization leads to a performance degradation ranging from $5\%$ to $70\%$.
9 Related Work
@@
Many existing frameworks, such as TensorFlow XLA [1, 9], PyTorch [34], and TensorRT [42], rely on GPU experts to manually design kernels for ML operators. Recently, significant engineering effort has been dedicated to hand-optimizing GPU kernels for commonly used DNNs, particularly foundation models [12]. For example, to accelerate attention computation [47], several specialized kernels have been developed based on FlashAttention [17, 5, 21, 4]. Due to the increasing complexity of modern GPUs—such as tensor cores in A100s [29] and thread block clusters in H100s [7]—manually designed kernels may miss subtle optimizations that are hard to discover manually.
@@
Superoptimization was originally introduced to find optimal instruction sequences [30, 36, 11]. Recent work has applied superoptimization techniques to tensor programs [25, 46, 52, 49, 45, 26, 23, 24]. However, all these attempts only consider algebraic transformations at the kernel level and cannot discover more sophisticated optimizations that require jointly considering algebraic and schedule transformations at all of the kernel, block, and thread levels. Our evaluation shows that Mirage largely outperforms existing DNN superoptimizers, demonstrating the importance of multi-level joint optimization.
@@
Recent work has introduced ML compilers that automatically optimize the execution schedule of kernel GPUs. Systems such as TVM [13, 14], Ansor [51], and Triton [43], along with others [53, 20, 19], build on the idea of algorithm-schedule separation introduced in Halide. They search for optimized schedules to execute a user-specified algorithm on GPUs. However, schedule-based approaches require users to explicitly specify the algorithm for each kernel, and their performance is limited to the quality of these provided algorithms.
@@
Welder [38] and ASPEN [33] introduce multi-level tile graphs that share similarities with Mirage’s $\mu$Graphs, as both representations follow the GPU hierarchy. However, prior work focuses on scheduling transformations, while Mirage extends beyond scheduling by also considering algebraic transformations and the discovery of new custom kernels. Most optimizations presented in this paper fall outside the scope of these prior approaches.
10 Conclusion
This paper proposes Mirage, the first multi-level superoptimizer for tensor programs. Mirage introduces a hierarchy graph representation to specify a tensor program at the kernel, thread block, and thread levels of the GPU execution hierarchy, and uses a novel pruning technique based on abstraction to significantly reduce the search space Mirage needs to consider while providing a certain optimality guarantee. Mirage outperforms existing tensor program optimizers by up to 3.3$\times$, even for widely used and heavily optimized DNNs.
Acknowledgment
We would like to thank the anonymous reviewers and our shepherd, Stephanie Wang, for their valuable comments and suggestions. We thank Tianqi Chen, Phillip Gibbons, Bohan Hou, Muyan Hu, Jinchen Jiang, Xiaoyu Jiang, Ruihang Lai, Yu Zhou, and other CMU Catalyst members for their feedback on this work. This research is partially supported by NSF awards CNS-2147909, CNS-2211882, and CNS-2239351, and research awards from Amazon, Cisco, Google, Meta, NVIDIA, Oracle, Qualcomm, and Samsung. This research is also partially supported by a research grant from the Center for New Scientists at the Weizmann Institute of Science and by a grant from the Azrieli Foundation.